OpenStudy (anonymous):

need help verifying please, still don't quite understand this one (sin/1 - cos) - (1 - cos/ sin) = 2cot It looks like I could cancel out the like terms but I'm not sure

OpenStudy (anonymous):

\[\frac{b}{1-a}-\frac{1-a}{b}\] is some algebra. do that first

OpenStudy (anonymous):

cross multiply first right?

OpenStudy (anonymous):

it is less confusing than working with sines and cosines. then put them back in the end i hope it is clear what i did , replace \(\sin(x)\) by \(b\) and \(\cos(x)\) by \(a\) yes, "cross multiply" although i think i would say "add"

OpenStudy (anonymous):

you will also need trig identities (sin^2(x) + cos^2(x) = 1

OpenStudy (anonymous):

just do the algebra, totally ignore the trig if you see \(a^2+b^2\) you can replace it by 1, but that is at the end oh, what @VeritasVosLiberabit said

OpenStudy (anonymous):

actually maybe you won't

OpenStudy (anonymous):

okay so it would then be sin^2/ 1 - cos2?

OpenStudy (anonymous):

find a common denominator

OpenStudy (anonymous):